If the perimeter of a rectangle is 18 ft and the width is 5 ft, find its length. [Hint: The perimeter of a rectangle is given by P = 2L + 2W.]
Found 2 solutions by sarah_adam, checkley71:
Answer by sarah_adam(201)

P = 2L + 2W
Given P = 18ft and W = 5ft
18 = 2L + 2(5)
18 = 2L +10
2L = 8
L = 8/2 = 4Ft

Answer by checkley71(8403)

18=2*5+2L OR 18=10+2L OR 8=2L OR L=4